OpenStudy (anonymous):

Help here please:) Determine an interest rate less than 15%

OpenStudy (experimentx):

is it a simple interest problem??

OpenStudy (experimentx):

the question is put up in a quite difficult way .. is it just asking to find the rate if you receive 500k for a million in two years??

OpenStudy (experimentx):

lol ... that what i hate about these type of questions. do you have answer?? if you have answer we might construct as well as simplify question

OpenStudy (experimentx):

damn silly question it says that t=? for A = 2P for r<15%

OpenStudy (experimentx):

let's put rate is 15 percent By the way, are you from english speaking country??

OpenStudy (experimentx):

I am not ... though it would seem you have more problem than me.

OpenStudy (experimentx):

lol ... nevermind. thats just a joke.

OpenStudy (experimentx):

it wouldn't be so wrong to put 15% just for now.

OpenStudy (experimentx):

it doesn't say anything about how many times it get's compounded in a year or it is saying, it get's compounded every two years.

OpenStudy (experimentx):

seems like i did not take regular deposit into account.

Directrix (directrix):

Where is it written that $500k has to be earned? That was an illustration of the meaning of the question, I thought. (***for example***, under what conditions will you have a future value of $1 000 000, having earned $500 000 interest?) But not the question.

OpenStudy (radar):

Agree, the question creates more questions.

OpenStudy (phi):

Use the FV formula here. \( FV= pmt \cdot \frac{(1+i)^n-1}{i} \) The total amount of money you put in is the payment amount times the number of payments: n*pmt We want the Future value to be twice what you put in: 2*n*pmt Plug this into the formula \( 2*n*pmt= pmt \cdot \frac{(1+i)^n-1}{i} \) or \( 2n=\frac{(1+i)^n-1}{i} \)

OpenStudy (phi):

To keep things simple, let's put money in once a year, and have the interest rate 10% per year. \[ 2n=\frac{(1.1)^n-1}{0.1} \]

OpenStudy (phi):

What are they asking: Determine an interest rate less than 15%, a period of investment greater than two years, and a regular payment that will result in the total amount of interest you earn being equal to the total amount of money you put in? They are asking you to pick a set of numbers (interest, payment size, investment time) so that at the end of the time period, you will have doubled your money. So I am going to assume 10% interest rate, and assume we put in a payment once a year. How long do we have to go to double our money?

OpenStudy (radar):

He was suggesting 10%, did you plug in 10% or 1%?

OpenStudy (radar):

I am here. was waiting for answer to my question.

OpenStudy (radar):

I see you have answered it.

OpenStudy (phi):

We could try making payments monthly to see if it matters.

OpenStudy (phi):

It that a regular payment or just a single deposit at the start?

OpenStudy (phi):

What is the total amount of money you put in?

OpenStudy (phi):

They want (for example, under what conditions will you have a future value of $1 000 000, having earned $500 000 interest?) So if the FV is 80000 they want the interest to be 40000

OpenStudy (phi):

Yes. Try 10% per year, and 14 years

OpenStudy (phi):

Yes, it is very very close

OpenStudy (phi):

I'm just saying 2*139749= 279498 which is close enough!

OpenStudy (phi):

To get the exact number of years we have to solve \( 0.2n+1 =(1.1)^n \) for n

OpenStudy (phi):

which I can do using Wolfram.

OpenStudy (phi):

Yes, especially if you make a note of the small difference.

OpenStudy (phi):

Here is a link to Wolfram's solution http://www.wolframalpha.com/input/?i=.2n%2B1+%3D+1.1%5En+ It is the 2nd red dot, about 14.02 years

OpenStudy (phi):

Time to go. Good luck with this stuff....